OpenStudy (anonymous):

5.7q-5.5-5.4q=.7q-3.7 I came up with q=4.5 but it does not check as right. Please help

OpenStudy (anonymous):

(5.7-5.4-0.7)q = (5.5-3.7) (-0.4)q=1.8 q=-4.5

OpenStudy (anonymous):

You missed the minus sign.

OpenStudy (anonymous):

hah yeah cakewalk is right, check your other post. I also posted the answer there =]